Find the equation of the line perpendicular to the line: 2x-3y=4 and passes through the point (-4,1)
Answer by robertb(5830)

The unknown line has the form 3x + 2y =c, where c is a constant. To find c, use the coordinates of the given point: 3*-4 + 2*1 = -12+ 2 = -10. Therefore the equation of the line is 3x + 2y = -10.